OpenStudy (anonymous):

Which polynomial is a difference of two squares? (4 points) x2 - 12 x2 + 12 x2 - 36 x2 + 36

OpenStudy (mathstudent55):

Is 12 a square or is 36 a square?

OpenStudy (shamil98):

Also the keyword there is "difference" of two squares

OpenStudy (anonymous):

36 @mathstudent55

OpenStudy (shamil98):

Good, now a difference of two squares means: (x-a)(x+a) = x^2 - ax + ax - a^2 = x^2 - a^2 which one of those are like this one?

OpenStudy (anonymous):

c

OpenStudy (mathstudent55):

Correct. You need 36. Also, "difference" means a subtraction.

OpenStudy (mathstudent55):

C. is correct.

OpenStudy (anonymous):

k thanxs

OpenStudy (mathstudent55):

You're welcome.